Hi can you please help me with some quadractic equations Thanks.(^2 means squared number)
x^2+7x+6=0
3x^2-5x-12=0
2x^2-5x-7=0
2.5x^2+6.2x=1.4

Answer by checkley75(3666)

X^2+7X+6=0, (X+6)(X+1)=0, X=-6 & X=-1
3X^2-5X-12=0, (3X+4)(X-3)=0, X=-4/3, X=3
THE REST I'LL LEAVE FOR YOU TO SOLVE!!!!!